# Information scrambling in quantum-walks

### v1

@@ II.1   Out-of-time-ordered correlators
III   Krylov complexity
III.1   Continuous-time evolution
@@
III.2   Discrete-time evolution
@@
To formulate the K-complexity for system with discrete-time evolution such that $\mathcal{O}_{t}=U^{\dagger}_{t}\mathcal{O}_{t-1}U_{t}$, where $t=1,2,\ldots$ and $U$ is the unitary operator describing system dynamics. We define the Krylov basis by choosing $|O_{0})=\mathcal{O}_{0}$ and then recursively orthogonalizing each $\mathcal{O}_{t}$ with all the $|O_{t})$ for $i<t$. At any $t$, we can expand the state $\mathcal{O}_{t}$ in Krylov basis as
$\mathcal{O}_{t}=\sum_{n=0}^{D}\phi_{n,t}|O_{n})\,,$ (7)
@@
where the expansion coefficient $\phi_{n,t}=(O_{n}|\mathcal{O}_{t})$. We define the K-complexity of the state as the average position of the distribution on the ordered Krylov basis:
$K(t)=\sum_{n=0}^{D}n|\phi_{n,t}|^{2}\,.$ (8)
IV   Discrete-time quantum walks
@@
$S=\lvert\downarrow\rangle\langle\downarrow\rvert\otimes T_{+}+\lvert\uparrow\rangle\langle\uparrow\rvert\otimes T_{-}\qquad\text{where}\quad T_{\pm}=\sum_{x\in\mathbb{Z}}|x\pm 1\rangle\langle x|$ (12)
@@
$v_{g}(k,\theta)\equiv\frac{d\omega(k,\theta))}{dk}=\pm\frac{k\cos\theta}{\sqrt{k^{2}\cos\theta+2(1-\cos\theta)}}\,.$ (15)
@@
IV.1   Disordered discrete-time quantum walk
@@
V   Results
V.1   Out-of-time-ordered correlator
@@
We consider the function $C_{\mu\nu}(l,t)$ in Eq. (19) for the discrete-time quantum walk in one-dimension and coin-angle $\theta$ for varying distance $l$ between the initial operators. Figure 2 shows the numerical results for $C_{\mu\nu}(l,t)$ at various time slices. We can identify the velocity of the wavefront as $v_{B}=\text{max}_{k}\ d\epsilon_{k}/dk=\text{max}_{k}\ v_{g}(k,\theta)$. In the present case, the OTOC function is “shell-like”. That is, inside the timelike region, in the long-time limit, $C_{\mu\nu}(l,t)\rightarrow 0$, indicating no scrambling of operator $W^{\mu}_{l}(t)$, the vanishing of the $C_{\mu\nu}$ OTOC in the long-time limit suggests that expansion of $W^{\mu}_{l}(t)$ in terms of Pauli strings does not contain many pauli matrices “in the middle” of the strings. The feature common to Integrable quantum systems [69].
@@
V.2   K-complexity
Considering the formulation of K-complexity presented in Sec. III for discrete-time evolution, we can show that the discrete-time quantum walk exhibit linear growth. We will consider the initial operator to be of the form $\mathcal{O}_{0}=\sigma^{\mu}\otimes|0\rangle\langle 0|$. At any time $t$, the operator is given by
@@ V.2   K-complexity
As the dispersion in discrete-time quantum walk grows linearly, the amplitude $\mathcal{A}_{0,2t}\approx 0$, and therefore, $(\mathcal{O}_{n}|\mathcal{O}_{m})\approx\delta_{nm}$. It follows that $\phi_{n,t}=(O_{n}|\mathcal{O}_{t})\approx\delta_{n,t}$, hence the K-complexity given by
$K(t)=\sum_{n}n|\phi_{n}(t)|^{2}\sim t\,.$ (27)
@@
In Fig. 5, the numerical analysis is presented for K-complexity of the discrete-time quantum walk. The norm of operator $|A_{n})$ saturates at a constant value close to $1$ showing that at late-time, the operators $|\mathcal{O}_{n})$ are approximately orthogonal to each-other. Therefore, the K-complexity obeys a linear growth.
@@
As we seen in Sec. V.1, the introduction of disorder leads to slowdown in information propagation. The K-complexity shows the similar localized behavior as OTOC in which it deviates from linear growth to suppress power law $\sim t^{1/\delta}$ with $\delta>1$. In this case, the amplitude $\mathcal{A}_{0,2n}$ goes to zero for large $n$ due to localization at $n=0$. Therefore, the expansion coefficient $\phi_{n,t}=(O_{n}|\mathcal{O}_{t})$ is non-zero for small $n$ which results in suppressed growth in complexity. In Fig. 6, the K-complexity is calculated under temporal and spatial disorder in discrete-time quantum walk for increasing disorder strength $W$. As with the OTOC, the suppression in K-complexity growth is smaller in temporal case compared to spatial as result of strong localization (or AL) in the later case.
VI   Conclusion

### v2

@@ II.1 Out-of-time-ordered correlators
II.2 Krylov complexity
II.2.1 Continuous-time evolution
@@
II.2.2 Discrete-time evolution
@@
To formulate the K-complexity for system with discrete-time evolution such that $\mathcal{O}_{n}=U^{\dagger}_{n}\mathcal{O}_{n-1}U_{n}=\mathcal{U}_{n}[\mathcal{O}_{n}]$, where $n=1,2,\ldots$, $U_{n}$ is the unitary operator which describes the evolution of the system at time-step $t=nT$ with step-size $T$, and $\mathcal{U}_{n}$ is unitary superoperator given by $\mathcal{U}_{n}=U_{n}^{\dagger}\bullet U_{n}$. We define the Krylov basis by choosing $|O_{0})=\mathcal{O}_{0}$ and then recursively orthogonalizing each $\mathcal{O}_{n}$ with all the $|O_{n})$ for $i<n$. This choice of basis (represented by $\mathbb{O}$) maximizes the cost function defined as [59]
$K_{\mathbb{B}}(t)=\sum_{i}\zeta_{i}\lvert(\mathcal{O}_{t}|B_{i})\rvert^{2}$ (7)
with respect to arbitrary choice of orthonormal set $\mathbb{B}=\{|B_{i}):i=0,1,2,\ldots\}$. The coefficients $\zeta_{i}$ (refer to as “weight function”) are positive, increasing sequence of real numbers. The optimal choice of the basis follows from induction method. We fix the initial state as the first state of Krylov basis $\mathcal{O}_{0}=|O_{0})$. Assume the first $N$ vectors of certain basis $\mathbb{B}$ are same as the Krylov basis, i.e., $|B_{i}\rangle=|O_{i}\rangle$ for $i=0,1,\ldots,N-1$. By assumption
$\mathcal{O}_{n}=\sum_{j=0}^{N-1}(K_{j}|\mathcal{O}_{n})|O_{j})\qquad n\leq N-1\,.$
Therefore, the cost associated with two basis are same for $n\leq N-1$. The next state $|O_{N}\rangle$ into a part belonging to the Krylov subspace $|O_{i}\rangle$, for $i\leq N-1$, and a part perpendicular to it i.e.
$\mathcal{O}_{N}=p_{\perp}|O_{N})+p_{\parallel}|O_{\parallel}),\qquad|O_{\parallel}\rangle=\sum_{i=0}^{N-1}a_{i}|O_{i})$
where $|O_{N})$ is the next element of the Krylov basis by definition. A basis different from the Krylov one would necessarily not include $|O_{N}\rangle$. Therefore, the cost at discrete time $N$ would be larger, since we would have to express $|K_{N}\rangle$ in the new basis, which would require at least two vectors. Since the contribution to the cost from the part $|O_{\parallel}\rangle$ is the same in both bases, the cost must increase when we divide $|O_{N}\rangle$ into several contributions, since $\zeta_{n}$ is a strictly increasing function of $n$. Therefore, the Krylov basis minimizes the cost function for all times.
At any time $t=nT$, we can expand the state $\mathcal{O}_{t}$ in the Krylov basis as
$\mathcal{O}_{n}=\sum_{i=0}^{D^{2}}\phi_{i,n}|O_{i})\,,$ (8)
@@
where the expansion coefficient $\phi_{i,t}=(O_{i}|\mathcal{O}_{n})$. We define the K-complexity of the state as the average position of the distribution on the ordered Krylov basis:
$K(t)=\sum_{i=0}^{D^{2}}i|\phi_{i,n}|^{2}\,.$ (9)
Since the evolution operator itself is used as generator of Krylov basis, we can bound the maximum possible growth of K-complexity. For any system evolving unitarily in discrete-time, the K-complexity can grow at most linearly with time $t$. To prove this, we consider maximizing the complexity at anytime $t$ with respect to expansion coefficients, i.e.,
$\max_{\{\lvert\phi_{i,n}\rvert\}}K(t)=\max_{\{\lvert\phi_{i,n}\rvert\}}\sum_{i}i|\phi_{i,n}|^{2}$ (10)
with constraint $\sum_{i}|\phi_{i,n}|^{2}=1$. From the orthonormalization construction, it follows that the expansion coefficients $\phi_{i,n}=0\ \forall\ i>n$. Therefore, it follows that the complexity can at most grows linearly with time $t$ corresponding to case where $\phi_{i,n}=\delta_{i,n}$. In terms of operator, it means that $(O_{n^{\prime}}|O_{n})\ \forall\ n^{\prime}<n$ which corresponds to maximally ergodic regime universal in chaotic systems [60].
In the case where the evolution operator is time-independent, we can define a quasi-Hamiltonian $H_{q}$ such that $U=e^{-iH_{q}T}$. In this case, the unitary operator $U$ takes upper Hessenberg form in Krylov basis [60], given by
$(O_{i}|U|O_{j})=\begin{cases}0&\text{if}\ j>i+1,\\ b_{j}&\text{if}\ j=i+1,\\ a_{j}c_{i}/c_{j}&\text{if}\ j<i+1,\end{cases}$ (11)
with $a_{0}=c_{0}$. The quasi-Hamiltonian $H_{q}$ can itself be used in continuous-time setting to define complexity using the notion discussed in Sec. II.2.1. To investigate the relation between the complexities, we consider the following expansion
$e^{i\mathcal{L}_{q}T}=\sum_{k=0}^{\infty}\frac{(iT)^{k}}{k!}\mathcal{L}_{q}^{k}\approx\sum_{k=0}^{K_{c}}\frac{(iT)^{k}}{k!}\mathcal{L}_{q}^{k}$ (12)
where $\mathcal{L}_{q}$ is Liouvillian superoperator associated with the Hamiltonian $H_{q}$. The Taylor series is truncated at order $K_{c}$ which depends on the operator $\mathcal{L}_{q}$, step-size $T$ and error tolerance [61]. From Equation (12), it follows that the operator at first time-step $\mathcal{O}_{1}$ will be linear combination of operators $\{\mathcal{L}^{k}_{q}\mathcal{O}_{0}\ |\ k=0,1,\ldots,K_{c}\}$. In other words, the first Krylov basis vector
$|O_{1})=\sum_{k=0}^{K_{c}}\psi_{k,1}|O^{(q)}_{k})$ (13)
where the set $\mathbb{O}^{(q)}=\{|O^{(q)}_{k})\ |\ k=0,1,\ldots\}$ corresponds to Krylov basis associated with Liouvillian $\mathcal{L}_{q}$. In general, we will assume that there exist cutoff value $K_{c}^{(i)}$ such that
$|O_{i})$ $=\sum_{k=0}^{K_{c}^{(i)}}\psi_{k,i}|O_{k}^{(q)})$ (14)
$\mathcal{O}_{n}$ $=\sum_{i=0}^{n}\phi_{i,n}|O_{i})=\sum_{k=0}^{K_{c}^{(n)}}\phi^{(q)}_{k,n}|O^{(q)}_{k})\,.$ (15)
Therefore, the K-complexity associated with the Liovillian $\mathcal{L}_{q}$ (refer to as quasi K-complexity) given by
$\begin{split}K^{(q)}(t)&=\sum_{k=0}^{K_{c}^{(n)}}k|\phi_{k,n}^{(q)}|^{2}=\sum_{k=0}^{K_{c}^{(n)}}k\left|\sum_{i=0}^{n}\phi_{i,n}\psi_{k,i}\right|^{2}\\ &\leq\sum_{k=0}^{K_{c}^{(n)}}k\sum_{i=0}^{n}|\phi_{i,n}|^{2}|\psi_{k,i}|^{2}\\ &=\sum_{i=0}^{n}\left(\sum_{k=0}^{K_{c}^{(n)}}k|\psi_{k,i}|^{2}\right)|\phi_{i,n}|^{2}\,.\end{split}$ (16)
Therefore, the quasi K-complexity $K^{(q)}(t)$ is similar to that of the unitary operator with a modified weight factor $\zeta_{i}$. The suitable choice of weight factor to recover the quasi K-complexity is non-trivial and depends on growth of wave functions $\phi^{(q)}_{k,n}$. Let us consider case where the wave function grows exponentially $|\phi^{(q)}_{k,t}|^{2}\sim e^{-k/\xi(t)}$, where $\xi(t)$ is delocalization length that grows exponentially in time $\xi(t)\sim e^{2\alpha t}$ for $\alpha t\gg 1$, corresponding to exponential growth (maximum possible growth) of quasi K-complexity [19] $K^{(q)}(t)\sim e^{2\alpha t}$. To retain this growth, one possible choice of weight function is $\zeta_{i}=e^{2\alpha i}$ along with maximal growth of wave function $\phi_{i,n}=\delta_{i,n}$.
We will now consider the limiting case where step-size $T\rightarrow 0$ to recover the continuous limit. In this case, $e^{i\mathcal{L}_{q}T}\approx I+iT\mathcal{L}_{q}$, up to correction of order $\mathcal{O}(T^{2})$, therefore, the krylov basis generated by unitary operator $U$ and quasi-Liouvillian $\mathcal{L}_{q}$ matches exactly to each other. However, it should be noted that the two complexities will eventually start to differ at large time-step as error start to accumulate and grow large.
III Discrete-time quantum walks
@@
$S=\lvert\downarrow\rangle\langle\downarrow\rvert\otimes T_{+}+\lvert\uparrow\rangle\langle\uparrow\rvert\otimes T_{-},\quad T_{\pm}=\sum_{x\in\mathbb{Z}}|x\pm 1\rangle\langle x|$ (20)
@@
$\begin{split}v_{g}(k,\theta)&\equiv\frac{d\omega(k,\theta)}{dk}\\ &=\pm\frac{k\cos\theta}{\sqrt{k^{2}\cos\theta+2(1-\cos\theta)}}\,.\end{split}$ (23)
@@
III.1 Disordered discrete-time quantum walk
@@
IV Results
IV.1 Out-of-time-ordered correlator
@@
We consider the function $C_{\mu\nu}(l,t)$ in Eq. (27) for the discrete-time quantum walk in one-dimension and coin-angle $\theta$ for varying distance $l$ between the initial operators. Figure 2 shows the numerical results for $C_{\mu\nu}(l,t)$ at various time slices. We can identify the velocity of the wavefront as $v_{B}=\text{max}_{k}\ d\epsilon_{k}/dk=\text{max}_{k}\ v_{g}(k,\theta)$. It follows that light cone grows linearly with maximum velocity $v_{B}=1$ corresponding to $\theta=0$. In continuous systems, the long range interaction can result in super-linear growth of light-cone [1]. Therefore, in principle, the long range connectivity in discrete-time evolution can also result in super-linear growth of light-cone. In the present case, the OTOC function is “shell-like”. That is, inside the timelike region, in the long-time limit, $C_{\mu\nu}(l,t)\rightarrow 0$, indicating no scrambling of operator $W^{\mu}_{l}(t)$, the vanishing of the $C_{\mu\nu}$ OTOC in the long-time limit suggests that expansion of $W^{\mu}_{l}(t)$ in terms of Pauli strings does not contain many pauli matrices “in the middle” of the strings. The feature common to Integrable quantum systems [72]. Additionally, in case of quantum walk, we find that the functioon $C_{\mu\nu}$ oscillates between a finite value and zero which stems from the form of the shift operator and the choice of initial operators.
In integrable systems such as quantum spin systems [72], one can analytically find the squared commutator to exactly describe the nature of the wavefront and decay. And, although, quantum walk is also an integrable model, it turns out to be particularly hard to find out exact form of squared commutator. This is because the initial operators are local in position space, and the inverse fourier transform becomes hard due to the complexity of functions. However, in case of the Hadamard quantum walk ($\theta=\pi/4$), we can use the known asymptotic results of wavefunction to vaguely argue the nature of decay and wavefront [73]. To this end, the asymptotics for the wave function can be described in three regions using $\alpha=l/t$. The wave function is essentially uniformly spread over the interval between $\mp 1/\sqrt{2}=\mp v_{B}$ where its gross behaviour is like $1/\sqrt{t}$. Outside this interval, the wave function dies out very faster than any inverse polynomial in $t$. At the “wavefront” (±$\pm v_{B}$), there are two peaks of width $\mathcal{O}(t^{1/3})$, where the wave function goes as $t^{-1/3}$. If we approximately translate these wave function behaviour to operator $W^{\mu}_{l}(t)$, we conclude that the squared commutator decays as $\sim 1/t^{2}$ in the long-time limit and goes as $t^{-4/3}$ at the wavefront.
@@
IV.2 K-complexity
Considering the formulation of K-complexity presented in Sec. II.2 for discrete-time evolution, we can show that the discrete-time quantum walk exhibit linear growth. We will consider the initial operator to be of the form $\mathcal{O}_{0}=\sigma^{\mu}\otimes|0\rangle\langle 0|$. At any time $t$, the operator is given by (we will fix the step-size $T$ to unity.)
@@ IV.2 K-complexity
As the dispersion in discrete-time quantum walk grows linearly [63], the operator support on a particular site decays over time. Hence, the overlap between the initial operator $\mathcal{O}_{0}$ and operator at any subsequent time $\mathcal{O}_{2t}$ also decays over time. In other words, $\mathcal{A}_{0,2t}\approx 0$, and therefore, $(\mathcal{O}_{n}|\mathcal{O}_{m})\approx\delta_{nm}$. It follows that $\phi_{n,t}=(O_{n}|\mathcal{O}_{t})\approx\delta_{n,t}$, hence the K-complexity given by
$K(t)=\sum_{n}n|\phi_{n}(t)|^{2}\sim t$ (35)
therefore, grows linearly. In Fig. 5, the numerical analysis is presented for K-complexity of the discrete-time quantum walk. The norm of operator $|A_{n})$ saturates at a constant value close to $1$ showing that at late-time, the operators $|\mathcal{O}_{n})$ are approximately orthogonal to each-other. Therefore, the K-complexity obeys a linear growth.
@@
As we seen in Sec. IV.1, the introduction of disorder leads to slowdown in information propagation. The K-complexity shows the similar localized behavior as OTOC in which it deviates from linear growth to suppress power law $\sim t^{1/\delta}$ with $\delta>1$. In this case, the amplitude $\mathcal{A}_{0,2n}$ goes to zero for large $n$ due to localization at $n=0$. Therefore, the expansion coefficient $\phi_{n,t}=(O_{n}|\mathcal{O}_{t})$ is non-zero for small $n$ which results in suppressed growth in complexity. In Fig. 6, the K-complexity is calculated under temporal and spatial disorder in discrete-time quantum walk for increasing disorder strength $W$. As with the OTOC, the suppression in K-complexity growth is smaller in temporal case compared to spatial as result of strong localization (or AL) in the later case.
V Conclusion